OpenStudy (unklerhaukus):

Use transform definitions, and the evaluation of a suitable double integral, to calculate the Laplace transform of the following integral \[\mathcal L\left\{\int\limits_0^tx(u)\cdot\text du\right\}\]

OpenStudy (unklerhaukus):

\[\begin{align*}\mathcal L\left\{\int\limits_0^tx(u)\cdot\text du\right\} &=\int\limits_0^\infty\left(\int\limits_0^tx(u)\cdot\text du\right)e^{-pt}\cdot\text dt\\ &=\\ &=\\ &=\end{align*}\]

OpenStudy (anonymous):

I'm guessing that this involve integration by parts...

OpenStudy (unklerhaukus):

do i have to reverse the order of integration ?

OpenStudy (anonymous):

Not quite sure though.

OpenStudy (unklerhaukus):

@Zarkon

OpenStudy (unklerhaukus):

\[\begin{align*}\mathcal L\left\{\int\limits_0^tx(u)\cdot\text du\right\} &=\int\limits_0^\infty\left(\int\limits_0^tx(u)\cdot\text du\right)e^{-pt}\cdot\text dt\\ &=\int\limits_0^\infty\int\limits_0^ux(u)e^{-pt}\cdot\text dt\cdot\text du\\ ?&=\int\limits_0^\infty x(u)\int\limits_0^ue^{-pt}\cdot\text dt\cdot\text du\\ \end{align*}\]

OpenStudy (unklerhaukus):

@mahmit2012

OpenStudy (anonymous):

Somehow u become a t

OpenStudy (unklerhaukus):